OpenStudy (anonymous):

HELP!!! Use the trapezoidal rule with n = 4 to approximate the area bounded by the curves y = sin2x, the lines x = 0, y = 0, and x = 1. I. 0.277 II. 0.273 III. 0.555 IV. 1.109 V. 2.219

OpenStudy (amistre64):

define the rule ...

OpenStudy (amistre64):

average heights times width of partition

OpenStudy (amistre64):

\[\frac{b-a}{n}\left(\frac{f(x_0)+f(x_1)}{2}+\frac{f(x_1)+f(x_2)}{2}+...+\frac{f(x_{n-1})+f(x_n)}{2}\right)\] \[\frac{b-a}{2n}\left(f(x_0)+f(x_1)+f(x_1)+f(x_2)+...+f(x_{n-1})+f(x_n)\right)\] \[\frac{b-a}{2n}\left(f(x_0)+2f(x_1)+2f(x_2)+...+2f(x_{n-1})+f(x_n)\right)\]

OpenStudy (anonymous):

Still a little confused! Sorry...

OpenStudy (amistre64):

find the values for the heights ... sin(2(0)) sin(2(.25)) | sin(2(.50)) | doubles these middle values sin(2(.75)) | sin(2(1)) and add em all up

OpenStudy (amistre64):

is that sin(2x) or sin^2 (x) ??

OpenStudy (anonymous):

sin^2(x)

OpenStudy (anonymous):

0.277 is the answer! Thank you so much...!

OpenStudy (amistre64):

the rule is defined in your material ... the hard part is keeping it all together tho

OpenStudy (amistre64):

good luck

OpenStudy (anonymous):

Thanks!